### v1

@@ 3.1 One eigenvalue problem
$\langle Sp,q\rangle$ $=b(\bm{u}_{h},q_{h})+s(p_{h},q_{h}),$
$b(\bm{v}_{h},p_{h})$ $=a(\bm{u}_{h},\bm{v}_{h})+a_{s}(\bm{u}_{h},\bm{v}_{h}),\quad\forall\bm{v}_{h}\in\bm{V}_{h}.$
@@
(15) $\langle Sp,q\rangle$ $\leq\|{\rm div}\bm{u}_{h}\|\|q_{h}\|+\frac{1}{\nu+\gamma}s(p_{h},p_{h})^{\frac{1}{2}}s(q_{h},q_{h})^{\frac{1}{2}},$
(16) $a(\bm{u}_{h},\bm{u}_{h})+\gamma\|{\rm div}\bm{u}_{h}\|^{2}$ $=b(\bm{u}_{h},p_{h})\leq\|{\rm div}\bm{u}_{h}\|\|p_{h}\|.$
Using $\|{\rm div}\bm{u}_{h}\|\leq\|\nabla\bm{u}_{h}\|$ and $a(\bm{u}_{h},\bm{u}_{h})=\nu\|\nabla\bm{u}_{h}\|^{2}$, we obtain from (16) that
$(\nu+\gamma)\|{\rm div}\bm{u}_{h}\|\leq\|p_{h}\|.$
@@
$\langle Sp,q\rangle$ $\leq\frac{1}{\nu+\gamma}\left[\|p_{h}\|\|q_{h}\|+s(p_{h},p_{h})^{\frac{1}{2}}s(q_{h},q_{h})^{\frac{1}{2}}\right]$
$\leq\|\frac{1}{\sqrt{\nu+\gamma}}p_{h}\|\|\frac{1}{\sqrt{\nu+\gamma}}q_{h}\|+\frac{1}{\nu+\gamma}s(p_{h},p_{h})^{\frac{1}{2}}s(q_{h},q_{h})^{\frac{1}{2}}$
$\leq\|\widehat{S}^{\frac{1}{2}}p\|_{\ell^{2}}\|\widehat{S}^{\frac{1}{2}}q\|_{\ell^{2}}.$
@@
(17) $\frac{\langle Sp,q\rangle}{\|\widehat{S}^{\frac{1}{2}}p\|_{\ell^{2}}\|\widehat{S}^{\frac{1}{2}}q\|_{\ell^{2}}}\leq 1,$
@@ 4 Block preconditioner and convergence analysis
$\nu(\nu+\gamma)\langle\widehat{S}p,p\rangle\leq\langle Sp,p\rangle.$
Setting $p=\widehat{S}^{-1}q$ in the above inequality leads to (21).
@@
$\langle S^{-1}q,q\rangle=\langle S^{-1}\hat{S}\hat{q},\widehat{S}\hat{q}\rangle=\langle p,\widehat{S}\hat{q}\rangle=(\gamma+\nu)^{-1}\left((p_{h},\hat{q}_{h})+s(p_{h},\hat{q}_{h})\right),$
where $p$ solves
$A_{\gamma}u+B^{T}p=0\quad\text{with}\quad-Bu+Cp=(\gamma+\nu)^{-1}Mq+Cq.$
@@
(23) $\nu(\nabla\bm{u}_{h},\nabla\bm{v}_{h})+\gamma({\rm div}\bm{u}_{h},{\rm div}\bm{v}_{h})+(\bm{a}\cdot\nabla\bm{u}_{h},\bm{v}_{h})$ $=(p_{h},{\rm div}\bm{v}_{h}),$
(24) $({\rm div}\bm{u}_{h},\tau_{h})+\frac{1}{\nu+\gamma}s(p_{h},\tau_{h})$ $=\frac{1}{\nu+\gamma}\left[(\hat{q}_{h},\tau_{h})+s(\hat{q}_{h},\tau_{h})\right],$
for all $\bm{v}_{h}\in\bm{V}_{h}$, $\tau_{h}\in Q_{h}$. Taking $\tau_{h}=p_{h}$ in (24) and $\bm{v}_{h}=\bm{u}_{h}$ in (23), we have
$\langle S^{-1}q,q\rangle=({\rm div}\bm{u}_{h},p_{h})+\frac{1}{\nu+\gamma}s(p_{h},p_{h})=\nu\|\nabla\bm{u}_{h}\|^{2}+\gamma\|{\rm div}\bm{u}_{h}\|^{2}+\frac{1}{\nu+\gamma}s(p_{h},p_{h}).$
@@
$\langle\widehat{S}\hat{q},\hat{q}\rangle$ $=({\rm div}\bm{u}_{h},\hat{q}_{h})+\tfrac{1}{\nu+\gamma}s(p_{h},\hat{q}_{h})$
$\leq\|{\rm div}\bm{u}_{h}\|\|\hat{q}_{h}\|+\tfrac{1}{\nu+\gamma}s(p_{h},p_{h})^{\frac{1}{2}}s(\hat{q}_{h},\hat{q}_{h})^{\frac{1}{2}}$
$\leq\left((\gamma+\nu)\|{\rm div}\bm{u}_{h}\|^{2}+\tfrac{1}{\nu+\gamma}s(p_{h},p_{h})\right)^{\frac{1}{2}}\tfrac{1}{\nu+\gamma}\left(\|\hat{q}_{h}\|^{2}+s(\hat{q}_{h},\hat{q}_{h})\right)^{\frac{1}{2}},$
$=\left((\gamma+\nu)\|{\rm div}\bm{u}_{h}\|^{2}+\tfrac{1}{\nu+\gamma}s(p_{h},p_{h})\right)^{\frac{1}{2}}\langle\widehat{S}\hat{q},\hat{q}\rangle^{\frac{1}{2}}.$
@@
$\begin{split}\langle\widehat{S}^{-1}q,q\rangle&=\langle\widehat{S}\hat{q},\hat{q}\rangle\leq(\gamma+\nu)\|{\rm div}\bm{u}_{h}\|^{2}+\frac{1}{\nu+\gamma}s(p_{h},p_{h})\leq\nu\|\nabla\bm{u}_{h}\|^{2}+\gamma\|{\rm div}\bm{u}_{h}\|^{2}+\frac{1}{\nu+\gamma}s(p_{h},p_{h})\\ &=\langle S^{-1}q,q\rangle,\end{split}$
@@
(25) $\|r_{k}\|_{\mathcal{M}^{-1}}\leq\left(1-\tilde{c}\nu(\nu+\gamma)\right)^{k/2}\|r_{0}\|_{\mathcal{M}^{-1}}=q^{k}\|r_{0}\|_{\mathcal{M}^{-1}},$
@@
For arbitrary $[v;p]\neq 0$, we need to show that
@@
To estimate the right-hand side in (26), one sets $p=\widehat{S}^{-1}q$ in the above quantity leading to
$(\lambda-1)^{2}$ $=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}B^{T}\widehat{S}^{-1}q,A_{\gamma}^{-T}B^{T}\widehat{S}^{-1}q\rangle}{\langle\widehat{S}^{-1}S\widehat{S}^{-1}q,q\rangle}=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}B^{T}p,A_{\gamma}^{-T}B^{T}p\rangle}{\langle Sp,p\rangle}$
$=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}B^{T}p,A_{\gamma}^{-T}B^{T}p\rangle}{\langle BA_{\gamma}^{-1}B^{T}p,p\rangle}=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}u,A_{\gamma}^{-T}u\rangle}{\langle A_{\gamma}^{-1}u,u\rangle}=\frac{1}{4}\frac{\langle A_{s}u_{1},u_{1}\rangle}{\langle A_{\gamma}u_{1},u_{1}\rangle}=\frac{1}{4}.$
To estimate the right-hand side in (27), one sets $p=S^{-1}q$ leading to
$(\lambda-1)^{2}=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}B^{T}S^{-1}q,A_{\gamma}^{-T}B^{T}S^{-1}q\rangle}{\langle S^{-1}q,q\rangle}=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}B^{T}p,A_{\gamma}^{-T}B^{T}p\rangle}{\langle Sp,p\rangle}=\frac{1}{4}.$
@@
$\alpha_{1}\langle\tilde{S}p,p\rangle\leq\langle\widehat{S}p,p\rangle\leq\alpha_{2}\langle\tilde{S}p,p\rangle,$
@@
From the above analysis, we see that the simple Schur complement preconditioner $\widehat{S}$ leads to a mesh-independent solver that is robust with respect to the critical parameter $\nu$ once $\gamma$ is sufficiently large. In $\mathcal{\widehat{P}}$, we leave open the question of how to approximate the (1,1) block. In the literature, several approaches based on direct LU/ILU factorization, the block upper-triangular part of $A_{\gamma}$, or the multigrid technique for the (1,1) block have been studied (see [2, 38, 6, 15]). Regarding field-of-values convergence analysis of GMRES with approximations to both $A_{\gamma}$ and $S$, similar results can be obtained by following the techniques in [28, 3]. However, this extension is beyond the scope of this work.
@@ 5 Numerical experiments
We consider two 2D flows to access the method accuracy and solver performance.
@@ 5.1 Driven cavity problem
Figure 3: Exponentially spaced streamlines of $Q_{2}-Q_{2}$ FE solutions without and with grad-div stabilization with $h=1/32$ for $Re=1000$.
Figure 4: Exponentially spaced streamlines of $Q_{2}-Q_{2}$ FE solutions without and with grad-div stabilization with $h=1/64$ for $Re=3200$.
@@
Figures 35 show cavity flow at Reynolds numbers 1000, 3200, and 5000, using log-uniformly distributed streamlines. The primal vortex and bottom secondary vortices are observed for $Re=1000$, with the appearance of a vortex near the upstream upper corner for $Re=3200$, and tertiary vortices becoming visible at $Re=5000$. This vortex dynamics pattern is typical and extensively documented in the literature. Once again, we recover very similar flow using $Q_{2}$ equal-order elements with $\gamma=0$ and $\gamma=0.1$.
@@ 5.2 Backward facing step problem
Figures 6 and 7 show streamlines and pressure of $Q_{2}$ solutions for $Re=150$ and $Re=800$. This time, solutions with and without grad-div stabilization are visually indistinguishable. The computed $Q_{1}$ solutions with $h=1/32$ and both $\gamma=0.1$ and $\gamma=0$ were visually similar to those shown in the figures.
@@
In Table 6, we present the iteration numbers for $Re=800$. We notice an increase in Picard’s iteration numbers compared to the $Re=150$ case, as expected. The reduction in GMRES iteration counts with $\gamma=0.1$ is even more pronounced in this scenario than for $Re=150$. Once again, the iteration count of GMRES with preconditioner $\mathcal{\widehat{P}}$ and $\gamma=0.1$ remains unaffected by variations in mesh parameters and elements, consistent with our convergence analysis.

### v2

@@ 3.1 One eigenvalue problem
$\langle S\varphi,q\rangle$ $=b(\bm{w}_{h},q_{h})+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{\nu+\gamma}}s(\varphi_{h},q_{h}),$
$b(\bm{v}_{h},\varphi_{h})$ $=a(\bm{w}_{h},\bm{v}_{h})+a_{s}(\bm{w}_{h},\bm{v}_{h}),\quad\forall\bm{v}_{h}\in\bm{V}_{h}.$
@@
(15) $\langle S\varphi,q\rangle$ $\leq\|{\rm div}\bm{w}_{h}\|\|q_{h}\|+\frac{1}{\nu+\gamma}s(\varphi_{h},\varphi_{h})^{\frac{1}{2}}s(q_{h},q_{h})^{\frac{1}{2}},$
(16) $a(\bm{w}_{h},\bm{w}_{h})+\gamma\|{\rm div}\bm{w}_{h}\|^{2}$ $=b(\bm{w}_{h},\varphi_{h})\leq\|{\rm div}\bm{w}_{h}\|\|\varphi_{h}\|.$
Using $\|{\rm div}\bm{w}_{h}\|\leq\|\nabla\bm{w}_{h}\|$ and $a(\bm{w}_{h},\bm{w}_{h})=\nu\|\nabla\bm{w}_{h}\|^{2}$, we obtain from (16) that
$(\nu+\gamma)\|{\rm div}\bm{w}_{h}\|\leq\|\varphi_{h}\|.$
@@
$\langle S\varphi,q\rangle$ $\leq\frac{1}{\nu+\gamma}\left[\|\varphi_{h}\|\|q_{h}\|+s(\varphi_{h},\varphi_{h})^{\frac{1}{2}}s(q_{h},q_{h})^{\frac{1}{2}}\right]$
$\leq\|\frac{1}{\sqrt{\nu+\gamma}}\varphi_{h}\|\|\frac{1}{\sqrt{\nu+\gamma}}q_{h}\|+\frac{1}{\nu+\gamma}s(\varphi_{h},\varphi_{h})^{\frac{1}{2}}s(q_{h},q_{h})^{\frac{1}{2}}$
$\leq\|\widehat{S}^{\frac{1}{2}}\varphi\|_{\ell^{2}}\|\widehat{S}^{\frac{1}{2}}q\|_{\ell^{2}}.$
@@
(17) $\frac{\langle S\varphi,q\rangle}{\|\widehat{S}^{\frac{1}{2}}\varphi\|_{\ell^{2}}\|\widehat{S}^{\frac{1}{2}}q\|_{\ell^{2}}}\leq 1,$
@@ 4 Block preconditioner and convergence analysis
$\nu(\nu+\gamma)\langle\widehat{S}g,g\rangle\leq\langle Sg,g\rangle.$
Setting $g=\widehat{S}^{-1}q$ in the above inequality leads to (21).
@@
$\langle S^{-1}q,q\rangle=\langle S^{-1}\hat{S}\hat{q},\widehat{S}\hat{q}\rangle=\langle g,\widehat{S}\hat{q}\rangle=(\gamma+\nu)^{-1}\left((g_{h},\hat{q}_{h})+s(g_{h},\hat{q}_{h})\right),$
where $g$ solves
$A_{\gamma}w+B^{T}g=0\quad\text{with}\quad-Bw+Cg=(\gamma+\nu)^{-1}M{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{q}+C\hat{q}}.$
@@
(23) $\nu(\nabla\bm{w}_{h},\nabla\bm{v}_{h})+\gamma({\rm div}\bm{w}_{h},{\rm div}\bm{v}_{h})+(\bm{a}\cdot\nabla\bm{w}_{h},\bm{v}_{h})$ $=(g_{h},{\rm div}\bm{v}_{h}),$
(24) $({\rm div}\bm{w}_{h},\tau_{h})+\frac{1}{\nu+\gamma}s(g_{h},\tau_{h})$ $=\frac{1}{\nu+\gamma}\left[(\hat{q}_{h},\tau_{h})+s(\hat{q}_{h},\tau_{h})\right],$
for all $\bm{v}_{h}\in\bm{V}_{h}$, $\tau_{h}\in Q_{h}$. Taking $\tau_{h}=g_{h}$ in (24) and $\bm{v}_{h}=\bm{w}_{h}$ in (23), we have
$\langle S^{-1}q,q\rangle=({\rm div}\bm{w}_{h},g_{h})+\frac{1}{\nu+\gamma}s(g_{h},g_{h})=\nu\|\nabla\bm{w}_{h}\|^{2}+\gamma\|{\rm div}\bm{w}_{h}\|^{2}+\frac{1}{\nu+\gamma}s(g_{h},g_{h}).$
@@
$\langle\widehat{S}\hat{q},\hat{q}\rangle$ $=({\rm div}\bm{w}_{h},\hat{q}_{h})+\tfrac{1}{\nu+\gamma}s(g_{h},\hat{q}_{h})$
$\leq\|{\rm div}\bm{w}_{h}\|\|\hat{q}_{h}\|+\tfrac{1}{\nu+\gamma}s(g_{h},g_{h})^{\frac{1}{2}}s(\hat{q}_{h},\hat{q}_{h})^{\frac{1}{2}}$
$\leq\left((\gamma+\nu)\|{\rm div}\bm{w}_{h}\|^{2}+\tfrac{1}{\nu+\gamma}s(g_{h},g_{h})\right)^{\frac{1}{2}}\tfrac{1}{\nu+\gamma}\left(\|\hat{q}_{h}\|^{2}+s(\hat{q}_{h},\hat{q}_{h})\right)^{\frac{1}{2}},$
$=\left((\gamma+\nu)\|{\rm div}\bm{w}_{h}\|^{2}+\tfrac{1}{\nu+\gamma}s(g_{h},g_{h})\right)^{\frac{1}{2}}\langle\widehat{S}\hat{q},\hat{q}\rangle^{\frac{1}{2}}.$
@@
$\begin{split}\langle\widehat{S}^{-1}q,q\rangle&=\langle\widehat{S}\hat{q},\hat{q}\rangle\leq(\gamma+\nu)\|{\rm div}\bm{w}_{h}\|^{2}+\frac{s(g_{h},g_{h})}{\nu+\gamma}\leq\nu\|\nabla\bm{w}_{h}\|^{2}+\gamma\|{\rm div}\bm{w}_{h}\|^{2}+\frac{1}{\nu+\gamma}s(g_{h},g_{h})\\ &=\langle S^{-1}q,q\rangle,\end{split}$
@@
(25) $\|r_{k}\|_{\mathcal{M}^{-1}}\leq\left(1-\tilde{c}\nu(\nu+\gamma)\right)^{k/2}\|r_{0}\|_{\mathcal{M}^{-1}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varrho^{k}}\|r_{0}\|_{\mathcal{M}^{-1}},$
@@
For arbitrary $[v;q]\neq 0$, we need to show that
@@
To estimate the right-hand side in (26), one sets $g=\widehat{S}^{-1}q$ in the above quantity leading to
$(\lambda-1)^{2}$ $=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}B^{T}\widehat{S}^{-1}q,A_{\gamma}^{-T}B^{T}\widehat{S}^{-1}q\rangle}{\langle\widehat{S}^{-1}S\widehat{S}^{-1}q,q\rangle}=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}B^{T}g,A_{\gamma}^{-T}B^{T}g\rangle}{\langle Sg,g\rangle}$
$=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}B^{T}g,A_{\gamma}^{-T}B^{T}g\rangle}{\langle BA_{\gamma}^{-1}B^{T}g,g\rangle}=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}w,A_{\gamma}^{-T}w\rangle}{\langle A_{\gamma}^{-1}w,w\rangle}=\frac{1}{4}\frac{\langle A_{s}w_{1},w_{1}\rangle}{\langle A_{\gamma}w_{1},w_{1}\rangle}=\frac{1}{4},$
where $w=B^{T}g$ and $w_{1}=A_{\gamma}^{-T}w$.
To estimate the right-hand side in (27), one sets $g=S^{-1}q$ leading to
$(\lambda-1)^{2}=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}B^{T}S^{-1}q,A_{\gamma}^{-T}B^{T}S^{-1}q\rangle}{\langle S^{-1}q,q\rangle}=\frac{1}{4}\frac{\langle A_{s}A_{\gamma}^{-T}B^{T}g,A_{\gamma}^{-T}B^{T}g\rangle}{\langle Sg,g\rangle}=\frac{1}{4}.$
@@
$\alpha_{1}\langle\tilde{S}q,q\rangle\leq\langle\widehat{S}q,q\rangle\leq\alpha_{2}\langle\tilde{S}q,q\rangle,$
@@
From the above analysis, we see that the simple Schur complement preconditioner $\widehat{S}$ leads to a mesh-independent solver that is robust with respect to the critical parameter $\nu$ once $\gamma$ is proportional to $\nu^{-1}$. In $\mathcal{\widehat{P}}$, we leave open the question of how to approximate the (1,1) block. In the literature, several approaches based on direct LU/ILU factorization, the block upper-triangular part of $A_{\gamma}$, or the multigrid technique for the (1,1) block have been studied (see [2, 38, 6, 15]). Regarding field-of-values convergence analysis of GMRES with approximations to both $A_{\gamma}$ and $S$, similar results can be obtained by following the techniques in [28, 3]. However, this extension is beyond the scope of this work.
@@ 5 Numerical experiments
We consider two 2D Navier–Stokes flows to access the method accuracy and solver performance.
@@ 5.1 Driven cavity problem
The computed vortex dynamics align with patterns extensively documented in the literature. For $Re=1000$, the primary vortex and bottom secondary vortices are observed, while a vortex near the upstream upper corner appears at $Re=3200$. At $Re=5000$, tertiary vortices become visible. Figure 3 illustrates the flow at a Reynolds number of 5000, showing a very similar pattern recovered using $Q_{2}$ equal-order elements with $\gamma=0$ and $\gamma=0.1$.
@@ 5.2 Backward facing step problem
Figures 4 and 5 show streamlines and pressure of $Q_{2}$ solutions for $Re=150$ and $Re=800$. This time, solutions with and without grad-div stabilization are visually indistinguishable. The computed $Q_{1}$ solutions with $h=1/32$ and both $\gamma=0.1$ and $\gamma=0$ were visually similar to those shown in the figures.
@@
In Table 6, we present the iteration numbers for $Re=800$. We notice an increase in Picard’s iteration numbers compared to the $Re=150$ case, as expected, as well as the slight increase in GMRES iterations. At the same time, the reduction in GMRES iteration counts with $\gamma=0.1$ is even more pronounced in this scenario than for $Re=150$. Once again, the iteration count of GMRES with preconditioner $\mathcal{\widehat{P}}$ and $\gamma=0.1$ remains unaffected by variations in mesh parameters and elements, consistent with our convergence analysis. The number of GMRES iterations per nonlinear step fluctuates very little from the averaged numbers reported in the tables.